### v3

@@ 3.2. Compatibly pointed elements and good elements
such that the set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has finitely many maximal elements with respect to the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. One can see that the collection $R_{t}$ has a ring structure.
@@
  • The set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has a unique maximal element ${\bf g}$ under the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. We denote by $\operatorname{\mathsf{deg}}^{t}(u):={\bf g}$ and call it the degree of $u$ with respect to the seed $({\bf x}_{t},\widetilde{B}_{t})$.
@@
By definition, each pointed element $u$ for the seed $({\bf x}_{t},\widetilde{B}_{t})$ takes the following form.
@@
    An element $u\in\mathcal{U}$ is said to be compatibly pointed, if $u$ is pointed for any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ and the collection $[{\bf g}]:=\{\operatorname{\mathsf{deg}}^{t}(u)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. In this case, we also say that $u$ is $[{\bf g}]$-pointed.
@@
    A $[{\bf g}]$-pointed element $u\in\mathcal{U}$ is said to be $[{\bf g}]$-good, if it is universally positive. In other words, the Laurent expansion of $u$ with respect to any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ can be written as
@@
    where ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is the value of $[{\bf g}]$ at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$.
Let $u$ be a $[{\bf g}]$-pointed element, then the Laurent expansion of $u$ with respect to any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ has a canonical expression:
(3.4) $u={\bf x}_{t}^{{\bf g}_{u}^{t}}F_{u}^{t}(\hat{y}_{1;t},\ldots,\hat{y}_{n;t}),$
where ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is the value of $[{\bf g}]$ at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
@@
  • For $k\in[1,n]$, we denote by $f_{k;u}^{t}$ the highest degree of $y_{k}$ in $F_{u}^{t}$. The vector
    ${\bf f}_{u}^{t}:=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}\in\mathbb{N}^{n}$
@@
  • The $[{\bf g}]$-pointed element $u$ is said to be $[{\bf g}]$-bipointed, if for any vertex $t\in\mathbb{T}_{n}$ the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in $F_{u}^{t}$ and it has coefficient $1$;
  • The $[{\bf g}]$-pointed element $u$ is said to be $[{\bf g}]$-bigood, if it is $[{\bf g}]$-bipointed and universally positive.
@@
Keep the setting as in Definition 3.6. Given a vertex $w\in\mathbb{T}_{n}$ and a vector ${\bf v}\in\mathbb{Z}^{m}$, we denote by $[({\bf v},w)]$ the unique tropical point $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ determined by the condition ${\bf g}^{w}={\bf v}$.
@@
Let ${\bf x}_{w}^{\bf v}$ be a cluster monomial in $\mathcal{U}$. Then ${\bf x}_{w}^{\bf v}$ is $[({\bf v},w)]$-bigood in $\mathcal{U}$.
@@
By the results in [GHKK18], we know that $u:={\bf x}_{w}^{\bf v}$ is $[({\bf v},w)]$-good in $\mathcal{U}$. Let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u={\bf x}_{w}^{\bf v}$ with respect to vertex $t\in\mathbb{T}_{n}$. Since the $F$-polynomials of cluster variables have constant term $1$ and by [fomin_zelevinsky_2007, Proposition 5.3], we know that the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in the $F$-polynomial $F_{u}^{t}$ with coefficient $1$. Hence, $u={\bf x}_{w}^{\bf v}$ is $[({\bf v},w)]$-bigood in $\mathcal{U}$. ∎
@@
with respect to the initial seed $({\bf x},\widetilde{B})$ are given as follows:
@@ 3.3. Basic properties of pointed elements
Lemma 3.9.
[qin_2019]*Lemmas 3.2.5, 3.2.6 Suppose that $u,u^{\prime}\in R_{t}$ are two pointed elements for seed $({\bf x}_{t},\widetilde{B}_{t})$. Then the following statements hold.
@@
    [GHKK18] Each column of $C_{t}^{w}$ is either a non-negative vector or a non-positive vector.
@@
    [NZ12, (3.11)] $SC_{t}^{w}S^{-1}(G_{t}^{w})^{\rm T}=I_{n}$, where $S$ is any skew-symmetrizer of $B_{w}$.
Lemma 3.11.
@@
    [qin_2019]*Lemma 3.3.6 Any Laurent monomial ${\bf x}_{t}^{\bf h}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and we have
@@
    [qin_2019]*Proposition 3.3.10 Let $\widehat{\bf y}_{t}$ and $\widehat{\bf y}_{w}$ be the collections of $\hat{y}$-variables at vertices $t$ and $w$. Then for any ${\bf v}\in\mathbb{Z}^{n}$, the Laurent monomial $\widehat{\bf y}_{t}^{\bf v}={\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and
@@
Notice that any element in $R_{w}$ can be written as a finite $\mathbb{Z}$-linear combinations of pointed elements for seed $({\bf x}_{w},\widetilde{B}_{w})$. By the above lemma, any Laurent monomial ${\bf x}_{t}^{\bf h}$ in $\mathcal{L}(t):=\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. Hence, $\mathcal{L}(t)$ is contained in $R_{w}$.
@@
Keep the above setting. Let $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$, and write ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ for $t\in\mathbb{T}_{n}$. Let ${w}$${s}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. The following statements hold.
@@
    Suppose that $u\in\mathcal{U}$ is a $[{\bf g}]$-pointed element. Let $u={\bf x}_{s}^{{\bf g}^{s}}\cdot\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ be the Laurent expansion of $u$ with respect to seed $({\bf x}_{s},\widetilde{B}_{s})$. Then the coefficient of $\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is $1$ and for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero.
@@
Because the $k$-th column of $C_{s}^{w}$ equals $-{\bf e}_{k}$, we have
@@
${\bf g}^{\prime}:=\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
@@
Using the equality $[b]_{+}g-[-b]_{+}g=bg=b[g]_{+}-b[-g]_{+}$ for any $b,g\in\mathbb{R}$, one can see the above equality is the same as
@@
(ii) Since $u$ is $[{\bf g}]$-pointed, we know that it is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. By Corollary 3.12 (ii), there exists a unique ${\bf u}\in\mathbb{N}^{n}$ with $c_{\bf u}=1$ such that
@@
Since the $k$th column of $C_{s}^{w}$ is $-{\bf e}_{k}$, we obtain that $C_{s}^{w}(d{\bf e}_{k}-[-g_{k}^{s}]_{+}{\bf e}_{k})$ belongs to $\mathbb{N}^{n}$ if and only if $d\leq[-g_{k}^{s}]_{+}$. Hence, for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero.
@@
Let $u$ be a $[{\bf g}]$-pointed element in $\mathcal{U}$ and let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u$ with respect to vertex $t\in\mathbb{T}_{n}$. The following statements hold.
@@
    Let $k$ be an integer in $[1,n]$ and $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$. Suppose that $f_{k;u}^{t}\neq 0$ but $f_{i;u}^{t}=0$ for any $i\in[1,n]\backslash\{k\}$. Then $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$.
@@
Since $u$ is $[{\bf g}]$-pointed in $\mathcal{U}$, the Laurent expansion of $u$ with respect to ${\bf x}_{t}$ has a cannonical expression:
@@
where ${\bf g}_{u}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ is the value of $[{\bf g}]$ at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
(i) By Corollary 3.13 (ii), we know that $y_{k}^{[-g_{k}^{t}]_{+}}$ must be a monomial appearing in $F_{u}^{t}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$. Recall that $f_{k;u}^{t}$ is the highest degree of $y_{k}$ in $F_{u}^{t}$. So we have $f_{k;u}^{t}\geq[-g_{k}^{t}]_{+}\geq 0$. Hence, if $f_{k;u}^{t}=0$, then $[-g_{k}^{t}]_{+}=0$. So $g_{k}^{t}$ is non-negative.
@@
(iii) Since $f_{i;u}^{t}=0$ holds for any $i\in[1,n]\setminus\{k\}$, we get that $F_{u}^{t}$ is a polynomial in the single variable $y_{k}$. In this case, Corollary 3.13 (ii) implies $f_{k;u}^{t}=[-g_{k}^{t}]_{+}$. Since $f_{k;u}^{t}\neq 0$, we have $g_{k}^{t}<0$ and thus $f_{k;u}^{t}=-g_{k}^{t}$.
Now $u$ has the following form:
@@
where $c_{d}\in\mathbb{Z}$. By the exchange relations, we have $x_{j;t^{\prime}}=x_{j;t}$ for any $j\neq k$ and $x_{k;t}\cdot x_{k;t^{\prime}}=M\cdot(1+\hat{y}_{k;t})$, where $M=\prod_{j=1}^{m}x_{j;t}^{[-b_{jk;t}]_{+}}$. By substituting these relations into the above equality and using the fact $u\in\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$, one deduces that
@@
and the Laurent expansion of $u$ with respect to seed $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})$ is a Laurent monomial in $\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$. Thus the $F$-polynomial $F_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is $1$. In particular, the $f$-vector ${\bf f}_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is zero. By (ii), we know that $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$.
4. $F$-invariant in cluster algebras
4.1. Maps between sets of tropical points
Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair. In this subsection, we give two maps. The first map is from the tropical points in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ and the second one is from the tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. The second map will play an important role in the proof Theorem 4.8.

### v4

@@ 3.2. Compatibly pointed elements and good elements
such that the set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has finitely many maximal elements with respect to the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. It is easy to see that the collection $R_{t}$ has a ring structure.
@@
  • The set $\{{\bf h}\in\mathbb{Z}^{m}\mid b_{\bf h}\neq 0\}$ has a unique maximal element $\bf g$} under the dominance order $\preceq_{t}$ on $\mathbb{Z}^{m}$. We denote by $\operatorname{\mathsf{deg}}^{t}(u)\coloneqq{\bf g}$ and call it the degree of $u$ with respect to the seed $({\bf x}_{t},\widetilde{B}_{t})$.
@@
By definition, each pointed element $u$ for the seed $({\bf x}_{t},\widetilde{B}_{t})$ takes the following form:
@@
    An element $u\in\mathcal{U}$ is said to be compatibly pointed, if $u$ is pointed for any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ and the collection ${\bf g}]\coloneqq\{\operatorname{\mathsf{deg}}^{t}(u)\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. In this case, we also say that $u$ is $\bf g$}]-pointed.
@@
    A $\bf g$}]-pointed element $u\in\mathcal{U}$ is said to be $\bf g$}]-good, if it is universally positive, i.e., $u\in\mathbb{Z}_{\geq 0}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ for any vertex $t\in\mathbb{T}_{n}$.
Let $u$ be a $\bf g$}]-pointed element, then the Laurent expansion of $u$ with respect to any seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{U}$ has a canonical expression:
@@
where ${\bf g}_{u}^{t}=\operatorname{\mathsf{deg}}^{t}(u)$ is the value of $\bf g$}] at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
@@
  • For $k\in[1,n]$, we denote by $f_{k;u}^{t}$ the maximal degree of $y_{k}$ in $F_{u}^{t}$. The vector
    ${\bf f}_{u}^{t}\coloneqq(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}\in\mathbb{N}^{n}$
@@
  • The $\bf g$}]-pointed element $u$ is said to be $\bf g$}]-bipointed, if for any vertex $t\in\mathbb{T}_{n}$ the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in $F_{u}^{t}$ and it has coefficient $1$;
  • The $\bf g$}]-pointed element $u$ is said to be $\bf g$}]-bigood, if it is $\bf g$}]-bipointed and universally positive.
@@
Keep the setting as in Definition 3.6. Given a vertex $w\in\mathbb{T}_{n}$ and a vector ${\bf v}\in\mathbb{Z}^{m}$, we denote by $({\bf v},w)]$ the unique tropical point $\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ determined by the condition ${\bf g}^{w}={\bf v}$.
@@
Let ${\bf x}_{w}^{\bf v}$ be a cluster monomial in $\mathcal{U}$. Then ${\bf x}_{w}^{\bf v}$ is $({\bf v},w)]-$bigood in $\mathcal{U}$.
@@
By the results in [GHKK18], we know that $u\coloneqq{\bf x}_{w}^{\bf v}$ is $({\bf v},w)]-$good in $\mathcal{U}$. Let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u={\bf x}_{w}^{\bf v}$ with respect to vertex $t\in\mathbb{T}_{n}$. Since the $F$-polynomials of cluster variables have constant term $1$ and by [fomin_zelevinsky_2007, Proposition 5.3], we know that the monomial $y_{1}^{f_{1;u}^{t}}\cdots y_{n}^{f_{n;u}^{t}}$ appears in the $F$-polynomial $F_{u}^{t}$ with coefficient $1$. Hence, $u={\bf x}_{w}^{\bf v}$ is $({\bf v},w)]-$bigood in $\mathcal{U}$. ∎
@@
[…]
@@ 3.3. Basic properties of pointed elements
Lemma 3.9 ([qin_2019]*Lemmas 3.2.5, 3.2.6).
Suppose that $u,u^{\prime}\in R_{t}$ are two pointed elements for seed $({\bf x}_{t},\widetilde{B}_{t})$. Then the following statements hold.
@@
    Each column of $C_{t}^{w}$ is either a non-negative vector or a non-positive vector.
@@
    $SC_{t}^{w}S^{-1}(G_{t}^{w})^{\rm T}=I_{n}$, where $S$ is any skew-symmetrizer of $B_{w}$.
Lemma 3.11 ([qin_2019]*Lem. 3.3.6, Prop. 3.3.10).
@@
    Any Laurent monomial ${\bf x}_{t}^{\bf h}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and we have
@@
    Let $\widehat{\bf y}_{t}$ and $\widehat{\bf y}_{w}$ be the collections of $\hat{y}-$variables at vertices $t$ and $w$. Then for any ${\bf v}\in\mathbb{Z}^{n}$, the Laurent monomial $\widehat{\bf y}_{t}^{\bf v}={\bf x}_{t}^{\widetilde{B}_{t}{\bf v}}$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$ and
@@
Notice that any element in $R_{w}$ can be written as a finite $\mathbb{Z}$-linear combinations of pointed elements for seed $({\bf x}_{w},\widetilde{B}_{w})$. By the above lemma, any Laurent monomial ${\bf x}_{t}^{\bf h}$ in $\mathcal{L}(t)\coloneqq\mathbb{Z}[x_{1;t}^{\pm 1},\ldots,x_{m;t}^{\pm 1}]$ is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. Hence, $\mathcal{L}(t)$ is contained in $R_{w}$.
@@
Keep the above setting. Let ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ be a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$, and write ${\bf g}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ for $t\in\mathbb{T}_{n}$. Let ${w}$${s}$$\scriptstyle{k}$ be an edge in $\mathbb{T}_{n}$. The following statements hold.
@@
    Suppose that $u\in\mathcal{U}$ is a $\bf g$}]-pointed element. Let $u={\bf x}_{s}^{{\bf g}^{s}}\cdot\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ be the Laurent expansion of $u$ with respect to seed $({\bf x}_{s},\widetilde{B}_{s})$. Then the coefficient of $\widehat{\bf y}_{s}^{[-g_{k}^{s}]_{+}{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is $1$ and for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero.
@@
Because the $k$th column of $C_{s}^{w}$ equals ${\bf e}_{k}$, we have
@@
${\bf g}^{\prime}\coloneqq\widetilde{G}_{s}^{w}{\bf g}^{s}-\widetilde{B}_{w}\cdot([-g_{k}^{s}]_{+}{\bf e}_{k}).$
@@
Using the equality $b]_{+}g-[-b]_{+}g=bg=b[g]_{+}-b[-g]_{+}$ for any $b,g\in\mathbb{R}$, one can see the above equality is the same as
@@
(ii) Since $u$ is $\bf g$}]-pointed, we know that it is pointed for the seed $({\bf x}_{w},\widetilde{B}_{w})$. By Corollary 3.12 (ii), there exists a unique ${\bf u}\in\mathbb{N}^{n}$ with $c_{\bf u}=1$ such that
@@
Since the $k$th column of $C_{s}^{w}$ is ${\bf e}_{k}$, we obtain that $C_{s}^{w}(d{\bf e}_{k}-[-g_{k}^{s}]_{+}{\bf e}_{k})$ belongs to $\mathbb{N}^{n}$ if and only if $d\leq[-g_{k}^{s}]_{+}$. Hence, for any integer $d$ with $d>[-g_{k}^{s}]_{+}$, the coefficient of $\widehat{\bf y}_{s}^{d{\bf e}_{k}}$ in $\sum\limits_{{\bf v}\in\mathbb{N}^{n}}c_{\bf v}\widehat{\bf y}_{s}^{\bf v}$ is zero. ∎
@@
Let $u$ be a $\bf g$}]-pointed element in $\mathcal{U}$ and let ${\bf f}_{u}^{t}=(f_{1;u}^{t},\ldots,f_{n;u}^{t})^{T}$ be the $f$-vector of $u$ with respect to vertex $t\in\mathbb{T}_{n}$. The following statements hold.
@@
[…]
@@
Since $u$ is $\bf g$}]-pointed in $\mathcal{U}$, the Laurent expansion of $u$ with respect to ${\bf x}_{t}$ has a canonical expression:
@@
where ${\bf g}_{u}^{t}=(g_{1}^{t},\ldots,g_{m}^{t})^{T}$ is the value of $\bf g$}] at vertex $t$ and $F_{u}^{t}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$.
(i) By Corollary 3.13 (ii), we know that $y_{k}^{[-g_{k}^{t}]_{+}}$ must be a monomial appearing in $F_{u}^{t}\in\mathbb{Z}[y_{1},\ldots,y_{n}]$. Recall that $f_{k;u}^{t}$ is the maximal degree of $y_{k}$ in $F_{u}^{t}$. So we have $f_{k;u}^{t}\geq[-g_{k}^{t}]_{+}\geq 0$. Hence, if $f_{k;u}^{t}=0$, then $g_{k}^{t}]_{+}=0$. So $g_{k}^{t}$ is non-negative.
@@
(iii) Since $f_{i;u}^{t}=0$ holds for any $i\in[1,n]\setminus\{k\}$, we get that $F_{u}^{t}$ is a polynomial in the single variable $y_{k}$. In this case, Corollary 3.13 (ii) implies $f_{k;u}^{t}=[-g_{k}^{t}]_{+}$. Since $f_{k;u}^{t}\neq 0$, we have $g_{k}^{t}<0$ and thus $f_{k;u}^{t}=-g_{k}^{t}$. Thus $u$ has the following form:
@@
where $c_{d}\in\mathbb{Z}$. By the exchange relations, we have $x_{j;t^{\prime}}=x_{j;t}$ for any $j\neq k$ and $x_{k;t}\cdot x_{k;t^{\prime}}=M\cdot(1+\hat{y}_{k;t})$, where $M=\prod_{j=1}^{m}x_{j;t}^{[-b_{jk;t}]_{+}}$. By substituting these relations into the above equality and using the fact $u\in\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$, we obtain that
@@
and the Laurent expansion of $u$ with respect to seed $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})$ is a Laurent monomial in $\mathbb{Z}[x_{1;t^{\prime}}^{\pm 1},\ldots,x_{m;t^{\prime}}^{\pm 1}]$. Thus the $F$-polynomial $F_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is $1$. In particular, the $f$-vector ${\bf f}_{u}^{t^{\prime}}$ of $u$ with respect to vertex $t^{\prime}$ is zero. By (ii), we know that $u$ is a cluster monomial in ${\bf x}_{t^{\prime}}$. ∎
4. Tropical invariant and $F$-invariant in cluster algebras
4.1. Semifield homomorphisms arising from good elements
Let $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$ be a Langlands-Poisson triple. In this subsection, we give a map from the tropical points in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$ to those in $\mathcal{S}_{X}(\mathbb{Z}^{\rm max})$. Thanks to this map, for each good element in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce a semifield homomorphism, which will be used to define tropical invariant and $F$-invariant.